# Audio-Visual Target Speaker Extraction with Reverse Selective Auditory Attention

### v1

@@ Abstract
Audio-visual target speaker extraction (AV-TSE) aims to extract the specific person’s speech from the audio mixture given auxiliary visual cues. Previous methods usually search for the target voice through speech-lip synchronization. However, this strategy mainly focuses on the existence of target speech, while ignoring the variations of the noise characteristics. That may result in extracting noisy signals from the incorrect sound source in challenging acoustic situations. To this end, we propose a novel reverse selective auditory attention mechanism, which can suppress interference speakers and non-speech signals to avoid incorrect speaker extraction. By estimating and utilizing the undesired noisy signal through this mechanism, we design an AV-TSE framework named Subtraction-and-ExtrAction network (SEANet) to suppress the noisy signals. We conduct abundant experiments by re-implementing three popular AV-TSE methods as the baselines and involving nine metrics for evaluation. The experimental results show that our proposed SEANet achieves state-of-the-art results and performs well for all five datasets. We will release the codes, the models and data logs.
@@ I Introduction
Figure 1: Typical approaches in AV-TSE focus on ‘Extraction’, which searches for the target speaker’s voice from the audio mixture to match the corresponding lip movements. However, the extraction results may involve noisy signals from incorrect sound sources. To alleviate these problems, we introduce the complementary ‘subtraction’ strategy. By analyzing the reverse selective auditory attention, our proposed method utilizes the estimated noisy signal and excludes them during extraction.
@@
The foundation of AV-TSE lies in the synchronization between lip movements and speech waveform [14, 15]. In an audio mixture, the target speech inside represents the pronunciation of phonemes, while the lip movements of the target speaker correspond to the viseme signals [16, 17] [^1] Viseme represents a basic visual speech unit that links to the phoneme of acoustic speech.. By establishing a temporal correlation between visemes and the corresponding speech signal components, it becomes possible to associate and align these two different modal signals [18]. Following this principle, an efficient AV-TSE solution is to encode visual lip frames into the embedding sequence, and feed it into the extractor module for obtaining the clean speech waveform. In particular, different architectures have been studied, such as convolutional neural networks (CNN) [4, 19, 20], temporal convolutional networks (TCN) [21, 14, 5, 22], cross-modal transformer [23, 6] and dual-path recurrent neural network networks (DPRNN) [24, 25].
@@
As depicted in the left part of Figure 1, most existing AV-TSE methods only focus on the ‘extraction’ strategy to search for the speech waveform component that is time-aligned with lip movements [5, 6, 4]. However, due to the generalization ability of the neural network, the extracted speech with a proper range of noise can also be considered synchronized with the lip movement by the model. So these methods may inadvertently involve noisy signals or even separate some voices from the wrong source [26]. That inspires us to guide the model in perceiving noisy signals during extraction. Additionally, some studies in image denoising and speech recognition have revealed the significance of investigating noisy signals as the auxiliary [27, 28, 29]. These findings motivate us to design a dedicated module that explicitly fits the AV-TSE task for exploring the speech-noise mutual exclusivity, i.e., reverse selective auditory attention. We summarise that as the ‘subtraction’ strategy (as indicated by the right part of Figure 1), which aims to estimate and understand the noisy signal in AV-TSE to improve training efficiency and system performance.
@@
In this paper, we propose a reverse attention mechanism to exclude the noisy information in AV-TSE to achieve this mutual exclusivity. It takes the target speech and the noisy signal for frame-level cross-attention. Suppose two audio frames from the respective signal achieve a high correlation score, these components should be penalized for incorrect extraction. This design can remove undesirable audio and avoid incorrect speaker extraction. Meanwhile, considering that noise includes interference speakers and background audio, which are all out-of-sync with the target speaker’s lip movements, we design a parallel speech and noise learning (PSNL) block to estimate speech and noisy signals at the same time. By inserting the reverse attention mechanism into the PSNL block, the model can perceive the noisy signal during the extraction process by combining the ‘extraction’ and ‘subtraction’ strategies.
@@
    Based on that, we propose the Subtraction-and-ExtrAction network (SEANet), a novel AV-TSE system with reverse selective auditory attention to assist the extraction by excluding the noisy signal. SEANet contains the PSNL block to estimate the speech and noisy signal.
@@ II-B1 Pre-enrolled speech
The widely-used auxiliary reference is the target speaker’s pre-enrolled speech utterance [36, 37]. Each speaker possesses the distinctive voice characterises, which can be modelled into a fixed dimensional vector, referred to as the speaker embedding [38, 39, 40]. Previous works [30, 35, 41] have employed the speaker encoder to derive speaker embedding from the reference speech, subsequently serving as attractors in the extraction process [42, 43]. Inspired by that, we propose to leverage the inherent consistency of the target speaker’s voice for AV-TSE. As illustrated by the blue lines in Figure 2, we expect that each frame of the extracted speech exhibits a high degree of correlation with the others due to the consistent voiceprint.
@@ II-B3 Speech-noise exclusivity
Noise always comes with speech, this correlation (speech-noise exclusivity) can provide complementary guidance for extraction. For example, in computer vision, extracting noise from images can contribute to the denoising process [45]; In speech recognition and enhancement, the consideration of noise can improve the system through a lightweight interaction layer [28, 29]. These findings demonstrate the importance of involving noise exploration. However, as illustrated by the red lines in Figure 2, this inter-frame mutual exclusivity between the target speaker’s speech and the noisy signal (refer as the ‘reverse selective auditory attention’ mechanism) has not been explored in AV-TSE. Meanwhile, previous studies in object detection and polyp segmentation [27, 46] studied the concept of reverse attention, which eliminates the core regions to make the model focus on local details. Inspired by that and modified based on the standard self-attention mechanism, we propose the reverse attention module. It penalizes segments with high speech-noise consistency since they should be mutually exclusive.
@@ III-A Task formulation and data generation
where $N$ is the number of interference speakers that $\geq 1$. Given the visual frames of the target speaker, which are denoted as $v$, AV-TSE aims to obtain $\hat{s}$ to approximate $s$. $v$ and $s$ should be synchronized.
As shown in Figure 3, in data generation process, we sum the speech from other speakers and the non-speech signal to generate the noisy signal: $n=\sum_{i=1}^{N}{o_{i}}+b$. Then $x$ can be obtained by summing $n$ and $s$. In our proposed SEANet, $v$ and $x$ are fed as the inputs which are encoded into the visual and audio embeddings to be fused and segmented into the multimodal embedding. In order to obtain the embedding of the target speaker’s speech and the noisy signal for the subsequent interaction, the pre-extractor and pre-suppressor modules are designed. Then several parallel speech and noise learning (PSNL) blocks are serially connected to learn the reverse selective auditory attention. Each PSNL block contains the reverse attention module to penalize the incorrect extraction. Finally, the embedding of the target speaker’s speech and the noisy signal are aggregated and decoded into the estimated target speaker’s speech $\hat{s}$ and the estimated noisy signal $\hat{n}$. Note that only $\hat{s}$ is required during evaluation.
@@ III-B Audio and visual encoder
Figure 3: The bottom panel is the data generation process for AV-TSE, the upper panel is our proposed SEANet. It extracts the clean speech of the target speaker from the mixture of audio. Specifically, SEANet contains $R$ repeated PSNL blocks to learn the reverse selective auditory attention between the estimated clean speech and noisy signal, $M_{si}$ and $M_{ni}$ are the output speech and noise embeddings from the $i^{th}$ block, respectively. $\oplus$ represents entrywise sum (to mix up audios) and $\otimes$ represents matmul product. Note that Pre-extractor and Pre-suppressor share the same model architecture with Extractor and Suppressor, respectively.
@@ III-C Audio-visual fusion
Previous research has confirmed that temporally aligning and concatenating audio and visual features is an effective approach for feature extraction [14, 5]. Based on that, the audio embedding $X$ is passed through a group normalization (GN) and a 1D-CNN layer. Subsequently, $X$ is concatenated with $V$ along the temporal dimension, which is then further refined by an additional 1D-CNN layer. The dimension of the integrated audio-visual embedding is $R^{D\times L}$ where $D$ represents the embedding dimension.
@@ III-D Pre-extractor and pre-suppressor
To achieve the subsequent speech-noise interaction process, we initially estimate the embeddings of the target speaker’s speech and the accompanying noise. As depicted in Figure 3, the audio-visual embedding $Y$ is fed into the pre-extractor and pre-suppressor to obtain the initial embedding $M_{s0}$ for the target speech and $M_{n0}$ for the noisy signal, respectively. These pre-processing modules are adaptations of DPRNN [24]. Note that the extractor and suppressor (which are elaborated in the subsequent PSNL section) share identical model architectures with their pre-processing counterparts. The details have been illustrated in Figure 4.
@@ III-E Parallel speech and noise learning (PSNL)
After that, parallel speech and noise learning (PSNL) block is used to learn the reverse selective auditory attention. $M_{s0}$ and $M_{n0}$ are fed into $R$ PSNL blocks. These blocks have non-shared parameter weights and are serially connected. In the $i^{th}$ PSNL block, the inputs are the speech embedding $M_{s(i-1)}$ and the noise embedding $M_{n(i-1)}$ from the $(i-1)^{th}$ PSNL block ($i\in[1,R]$). First, the intra-chunk attention (intra-att) module is used to learn the speech-noise mutual exclusivity within each chunk, which can avoid incorrect extraction in local frames. The output embeddings go through the linear layers with GN operation. Then, with a similar structure, an inter-att module is applied to compare the speech and noisy signals from different chunks to learn the global correctness. Here, the length of each chunk $K$ and the number of chunks $P$ are viewed as the sequence length in the intra-att and inter-att module, respectively. As shown in Figure 4, after the speech-noise interactor, an extractor and a suppressor are used to optimize the embedding of the target speaker’s speech and the noisy signal. This extractor (suppressor) have the same model architecture to the pre-extractor (pre-suppressor).
@@ III-F Reverse attention module
For reverse attention, the extracted speech should exclude the interference signal with low correlation: After multiplying the speech key and the noisy query, the output attention matrix goes through a reverse operation for the opposite attention matrix $A_{s-}$. Based on that, the high-corrected parts will be transferred into a low negative score to represent the incorrect extraction. The subsequent softmax operation $\sigma(\cdot)$ can ensure the final score matrix is in an appropriate range. The final score matrix $A_{s}$ is defined as:
$A_{s}=\frac{1}{2}(\underbrace{\sigma{(\frac{Q_{s}\cdot{K_{s}}^{T}}{\sqrt{D}}}}_{A_{s+}})+\underbrace{\sigma{(-\frac{Q^{\prime}_{n}\cdot{K_{s}}^{T}}{\sqrt{D}})}}_{A_{s-}})$ (2)
@@ III-G Audio decoder and constraints design
The final output embeddings are represented as $M_{sR}$ and $M_{nR}$. They are element-wise multiplied with $X$, respectively, followed by two audio decoder (shared weights) to obtain the estimated temporal target speech $\hat{s}_{R}\in R^{1\times T}$ and noisy signal $\hat{n}_{R}\in R^{1\times T}$. This decoder comprises a linear layer and an overlap-and-add (OnA) operation from signal processing.
@@ III-H Loss function
IV Experimental setting
IV-A Training set
IV-A1 LRS2

### v3

# Audio-Visual Target Speaker Extraction with Selective Auditory Attention

@@ Abstract
Audio-visual target speaker extraction (AV-TSE) aims to extract the specific person’s speech from the audio mixture given auxiliary visual cues. Previous methods usually search for the target voice through speech-lip synchronization. However, this strategy mainly focuses on the existence of target speech, while ignoring the variations of the noise characteristics, i.e., interference speaker and the background noise. That may result in extracting noisy signals from the incorrect sound source in challenging acoustic situations. To this end, we propose a novel selective auditory attention mechanism, which can suppress interference speakers and non-speech signals to avoid incorrect speaker extraction. By estimating and utilizing the undesired noisy signal through this mechanism, we design an AV-TSE framework named Subtraction-and-ExtrAction network (SEANet) to suppress the noisy signals. We conduct abundant experiments by re-implementing three popular AV-TSE methods as the baselines and involving nine metrics for evaluation. The experimental results show that our proposed SEANet achieves state-of-the-art results and performs well for all five datasets. The code can be found in: https://github.com/TaoRuijie/SEANet.git
@@ I Introduction
Figure 1: Typical approaches in AV-TSE focus on ‘Extraction’, which searches for the target speaker’s voice from the audio mixture to match the corresponding lip movements. However, the extraction results may involve noisy signals from incorrect sound sources. To alleviate these problems, we introduce the complementary ‘subtraction’ strategy. By analyzing this selective auditory attention, our proposed method utilizes the estimated noisy signal and excludes them during extraction.
@@
The foundation of AV-TSE lies in the synchronization between lip movements and speech waveform [14, 15]. In an audio mixture, the target speech inside represents the pronunciation of phonemes, while the lip movements of the target speaker correspond to the viseme signals [16, 17] [^1] Viseme: a basic visual speech unit links to the phoneme.. By establishing a temporal correlation between visemes and the corresponding speech signal components, it becomes possible to associate and align these two different modal signals [18]. Following this principle, an efficient AV-TSE solution is to encode visual lip frames into the embedding sequence, and feed it into the extractor module for obtaining the clean speech waveform. In particular, different architectures have been studied, such as convolutional neural networks (CNN) [4, 19, 20], temporal convolutional networks (TCN) [21, 14, 5, 22], cross-modal transformer [23, 6] and dual-path recurrent neural network (DPRNN) [24, 25].
@@
To achieve robust extraction, AV-TSE model needs to search for the clean target speech (synchronized to lip movements) and remove the noisy signal (unsynchronized to lip movements). However, based on our previous study of speech-lip synchronization [10], a well-trained neural network can tolerate the noisy signal during synchronization detection. As depicted in the left part of Figure 1, most existing AV-TSE methods are designed according on the ‘extraction’ strategy only. The emphasis is to guarantee that the output contains the speech waveform that be time-aligned with lip movements [5, 6, 4]. Based on that, AV-TSE models may inadvertently involve noisy signals or even separate some voices from the wrong source [26]. That inspires us to guide the model in perceiving noisy signals during extraction. Studies of negative prompting [27, 28] can guide the generation by providing unwanted signals. However, the noisy signals are unknown in AV-TSE, and the speech-noisy correlation is a complex temporal correlation. Additionally, some studies in image denoising and speech recognition have revealed the significance of investigating noisy signals as the auxiliary [29, 30, 31]. These motivate us to design a dedicated module that explicitly fits the AV-TSE task for exploring the speech-noise mutual exclusivity, i.e., selective auditory attention. We summarise that as the ‘subtraction’ strategy (as indicated by the right part of Figure 1), which aims to estimate and understand the noisy signal in AV-TSE to improve training efficiency and system performance.
@@
In this paper, we propose a reverse attention mechanism to exclude noisy information in AV-TSE to achieve this mutual exclusivity. It takes the target speech and the noisy signal for frame-level cross-attention. Suppose two audio frames from the respective signal achieve a high correlation score, these components should be penalized for incorrect extraction. This design can remove undesirable audio and avoid incorrect speaker extraction. Meanwhile, considering that noise includes interference speakers and background audio, which are all out-of-sync with the target speaker’s lip movements, we design a parallel speech and noise learning (PSNL) block to estimate speech and noisy signals at the same time. Furthermore, to extend this idea to the multi-modal processing, we design a multi-modal temporal attention module with diverse attempts to explore the effect of audio-visual correlation. By leveraging these modules, the framework can perceive the noisy signal during the extraction process by combining the ‘extraction’ and ‘subtraction’ strategies.
@@
    Based on that, we propose the Subtraction-and-ExtrAction network (SEANet), a novel AV-TSE system with selective auditory attention to assist the extraction by excluding the noisy signal. SEANet contains the PSNL block to estimate the speech and noisy signal. We also design the SEANet+ framework with the multi-modal temporal attention to explore the effect of visual information.
@@ II-B1 Pre-enrolled speech
The widely-used auxiliary reference is the target speaker’s pre-enrolled speech utterance [38, 39]. Each speaker possesses the distinctive voice characterizes, which can be modeled into a fixed dimensional vector, referred to as the speaker embedding [40, 41, 42]. Previous works [32, 37, 43] have employed the speaker encoder to derive speaker embedding from the reference speech, subsequently serving as attractors in the extraction process [44, 45]. Inspired by that, we propose to leverage the inherent consistency of the target speaker’s voice for AV-TSE. As illustrated by the blue lines in Figure 2, we expect that each frame of the extracted speech exhibits a high degree of correlation with the others due to the consistent voiceprint.
@@ II-B3 Speech-noise exclusivity
Noise always comes with speech, this correlation (speech-noise exclusivity) can provide complementary guidance for extraction. For example, in computer vision, extracting noise from images can contribute to the denoising process [47]; In speech recognition and enhancement, the consideration of noise can improve the system through a lightweight interaction layer [30, 31]. These findings demonstrate the importance of involving noise exploration. However, as illustrated by the red lines in Figure 2, this inter-frame mutual exclusivity between the target speaker’s speech and the noisy signal (refer as the ‘selective auditory attention’ mechanism) has not been explored in AV-TSE.
II-B4 Utilization of the negative singal
Utilizing the negative signal is verified to be efficient in multiple tasks. For instance, in text-to-image generation, the negative prompting solution [27, 28] can guide the generation process by providing the prompt to describe the unwanted signal. Meanwhile, previous studies [29, 48] in object detection and polyp segmentation 222the process of delineating the boundary of a polyp from the colon wall studied the concept of reverse attention, which eliminates the core regions to make the model focus on local details. However, compared with these non-temporal relationship, the speech-interference correlation is a more complex temporal relationship with the unknown interference presentation. Inspired by that and modified based on the standard self-attention mechanism, we propose the reverse attention module. It penalizes segments with high speech-noise and speech-vision consistency since they should be mutually exclusive.
@@ III-A Task formulation and data generation
where $N$ is the number of interference speakers that $\geq 1$. As shown in Figure 3, in data generation process, we sum the speech from other speakers and the non-speech signal to generate the noisy signal: $n=\sum_{i=1}^{N}{o_{i}}+b$. Then $x$ can be obtained by summing $n$ and $s$. Given the visual frames of the target speaker, which are denoted as $v$, AV-TSE aims to remove all noisy signal $n$ from $x$ and obtain $\hat{s}$ to approximate $s$. $v$ and $s$ should be synchronized. In our proposed SEANet, $v$ and $x$ are fed as the inputs which are encoded into the visual and audio embeddings to be fused and segmented into the multimodal embedding. In order to obtain the embedding of the target speaker’s speech and the noisy signal for the subsequent interaction, the pre-extractor and pre-suppressor modules are designed. Then several parallel speech and noise learning (PSNL) blocks are serially connected to learn the selective auditory attention. Each PSNL block contains the reverse attention module to penalize the incorrect extraction. Finally, the embedding of the target speaker’s speech and the noisy signal are aggregated and decoded into the estimated target speaker’s speech $\hat{s}$ and the estimated noisy signal $\hat{n}$. Note that only $\hat{s}$ is required during evaluation.
@@ III-B Audio and visual encoder
Figure 3: The bottom panel is the data generation process for AV-TSE, the upper panel is our proposed SEANet. It extracts the clean speech of the target speaker from the mixture of audio. Specifically, SEANet contains $R$ repeated PSNL blocks to learn the selective auditory attention between the estimated clean speech and noisy signal, $M_{si}$ and $M_{ni}$ are the output speech and noise embeddings from the $i^{th}$ block, respectively. $\oplus$ represents entrywise sum (to mix up audios) and $\otimes$ represents matmul product. Note that Pre-extractor and Pre-suppressor share the same model architecture with Extractor and Suppressor, respectively.
@@ III-C Audio-visual fusion
Previous research has confirmed that temporally aligning and concatenating audio and visual features is an effective approach for feature extraction [14, 5]. Based on that, the audio embedding $X$ is passed through a group normalization (GN) and a 1D-CNN layer. Subsequently, this embedding is concatenated with visual embedding $V$ along the temporal dimension, which is then further refined by an additional 1D-CNN layer. The dimension of the integrated audio-visual embedding is $R^{D\times L}$ where $D$ represents the embedding dimension, and $L$ is the number of audio frames, respectively.
@@ III-D Pre-extractor and pre-suppressor
To achieve the subsequent speech-noise interaction process, we initially estimate the embeddings of the target speaker’s speech and the accompanying noise. As depicted in Figure 3, the audio-visual embedding $Y$ is fed into the pre-extractor and pre-suppressor to obtain the initial embedding $M_{s0}$ for the target speech and $M_{n0}$ for the noisy signal, respectively. These pre-processing modules are adaptations of DPRNN [24]. Note that the extractor and suppressor (which are elaborated in the subsequent PSNL section) share identical model architectures with their pre-processing counterparts. More specifically, the extractor in Figure 4 has the same architecture as the pre-extractor, and the suppressor in Figure 4 has the same architecture as the pre-suppressor.
@@ III-E Parallel speech and noise learning (PSNL)
After that, parallel speech and noise learning (PSNL) block is used to learn the selective auditory attention. $M_{s0}$ and $M_{n0}$ are fed into $R$ PSNL blocks. These blocks have non-shared parameter weights and are serially connected. In the $i^{th}$ PSNL block, the inputs are the speech embedding $M_{s(i-1)}$ and the noise embedding $M_{n(i-1)}$ from the $(i-1)^{th}$ PSNL block ($i\in[1,R]$). First, the intra-chunk attention (intra-att) module is used to learn the speech-noise mutual exclusivity within each chunk, which can avoid incorrect extraction in local frames. The output embeddings go through the linear layers with GN operation. Then, with a similar structure, an inter-att module is applied to compare the speech and noisy signals from different chunks to learn the global correctness. Here, the length of each chunk $K$ and the number of chunks $P$ are viewed as the sequence length in the intra-att and inter-att module, respectively. As shown in Figure 4, after the speech-noise interactor, an extractor and a suppressor are used to optimize the embedding of the target speaker’s speech and the noisy signal. This extractor (suppressor) have the same model architecture to the pre-extractor (pre-suppressor).
@@ III-F Reverse attention module
For reverse attention, the extracted speech should exclude the interference signal with low correlation: After multiplying the speech key and the noisy query, the output attention matrix goes through a reverse operation for the opposite attention matrix $A_{s-}$. Based on that, the high-corrected parts will be transferred into a low negative score to represent the incorrect extraction. The subsequent softmax operation $softmax(\cdot)$ can ensure that the final score matrix is in a suitable range. The final score matrix $A_{s}$ is defined as:
$A_{s}=\frac{1}{2}(\underbrace{softmax{(\frac{Q_{s}\cdot{K_{s}}^{T}}{\sqrt{D}}}}_{A_{s+}})+\underbrace{softmax{(-\frac{Q^{\prime}_{n}\cdot{K_{s}}^{T}}{\sqrt{D}})}}_{A_{s-}})$ (2)
@@ III-G Audio decoder and constraints design
The final output embeddings are represented as $M_{sR}$ and $M_{nR}$. They are element-wise multiplied with $X$, respectively, followed by two audio decoder (shared weights) to obtain the estimated temporal target speech $\hat{s}_{R}\in R^{1\times T}$ and noisy signal $\hat{n}_{R}\in R^{1\times T}$. This decoder comprises a linear layer and an overlap-and-add (OLA) operation from signal processing.
@@ III-H Loss function
IV Exploration of multi-modal interaction
Figure 6: Multi-modal temporal attention block, which learns the interaction between the lip movements and the audio signal. F-SEANet system applies this module for audio-visual fusion, P-SEANet system inserts this module into the PSNL block after speech-noise interactor.
Compared with the audio-only TSE, AV-TSE has the advantages of leveraging the audio-visual correlation. Motivated by that, we adapt our design and propose different multi-modal interaction methods to boost SEANet. This utilizes lip movements to assist the target speech extraction and the interference signals suppression. Three different systems are studied: F-SEANet applies multi-modal temporal attention for the audio-visual fusion step, P-SEANet inserts the multi-modal temporal attention into the PSNE block, after the speech-noise interactor; A-SEANet adds the additional contrastive loss during training.
IV-A Fusion-based multi-modal interaction
The audio-visual fusion block in SEANet concentrates the obtained signal modality signal along the temporal domain. As shown in Fig 6, we design a multi-modal temporal attention to further explore our proposed reverse attention mechanism to explore more efficient audio-visual fusion.
Considering that the audio signal contains the target speech and the interference signal, in our multi-modal temporal attention, the visual signal interacts with the audio signal based on two branches: In both branches, the visual signal provides the query ($Q_{vs}$ and $Q_{vn}$), and the audio signal provides the key and value ($K_{s}$, $K_{n}$, $V_{s}$ and $V_{n}$). Here $F_{s}==F_{n}$, the difference is that one branch learns the positive audio-visual correlation to obtain the speech signal, while the other studies the negative audio-visual correlation and outputs the noisy signal. After that, these two signals ($F^{\prime}_{s}$ and $F^{\prime}_{n}$) are combined to get a robust audio representation. We name this system as F-SEANet.
IV-B PSNL-based multi-modal interaction
Then, we study the effect of conducting the multi-modal interaction into the PSNL block and name this system P-SEANet. First, to adjust the visual signal for multi-modal alignment, the output from the video encoder is segmented into several chunks to provide the visual query. After that, the multi-modal temporal attention module is inserted in each PSNL module, after each speech-noise interactor. This module contains architecture similar to the aforementioned one in the fusion block. The difference is that the inputs are the speech feature $F_{s}$ and the noisy feature $F_{n}$, and the outputs are the attentive speech and noisy feature ($F^{\prime}_{s}$ and $F^{\prime}_{n}$) without fusion.
IV-C Algorithm-based multi-modal interaction
Finally, we design the method to explore the multi-modal interaction from the algorithm aspect. Since the learned target speech has a positive correlation with the lip movements and the interference signal has a negative correlation, we design a contrastive learning based algorithm to involve the additional loss function and name it A-SEANet.
In A-SEANet, after each PSNL block, we first compute the cosine similarity between the speech and visual features along the time domain. Based on synchronization, these scores should be as high as possible; by contrast, the cosine similarity between the noisy and visual features ought to be as low as possible. These two similarity scores are subtracted as the additional training loss to update the system. We hope this can guide the framework in learning multi-modal temporal interaction.
V Experimental setting
V-A Training set
V-A1 LRS2